# Sliced Rényi Pufferfish Privacy: Directional Additive Noise Mechanism and Private Learning with Gradient Clipping

### v2

@@ Appendix D Minimality and Attainability of HUC
Fix $t$, a slice $u_{i}\in\mathbb{S}^{d-1}$, a prior $\theta\in\Theta$, and a secret pair $(s_{i},s_{j})\in\mathcal{Q}$ with $P^{S}_{\theta}(s_{i}),P^{S}_{\theta}(s_{j})>0$. Let $\gamma\in\Pi(\mu^{\theta}_{s_{i}},\mu^{\theta}_{s_{j}})$ be any admissible coupling of the dataset beliefs, and draw
$((X,X^{\prime}),R_{t})\sim\gamma\times\mathbb{P}_{\eta,\rho}.$
Define the update difference
$\widetilde{\Delta}_{t}(X,X^{\prime};R_{t}):=f_{t}(X,y_{<t};R_{t})-f_{t}(X^{\prime},y_{<t};R_{t}).$ (37)
By the HUC condition for $h_{t}$, we have
$\bigl|\langle\Delta_{t}(X,X^{\prime};R_{t}),u_{i}\rangle\bigr|\leq\sqrt{h_{t,i}}\quad\text{for }\gamma\times\mathbb{P}_{\eta,\rho}\text{-a.e. }((X,X^{\prime}),R_{t}).$ (38)
Step 1. Project onto the slice $u_{i}$ and write
$\widetilde{Y}_{t}:=\langle Y_{t},u_{i}\rangle=\langle f_{t}(X,y_{<t};R_{t}),u_{i}\rangle+\langle N_{t},u_{i}\rangle.$
Because $N_{t}\sim\mathcal{N}(0,\Sigma_{t})$ is independent of $((X,X^{\prime}),R_{t})$, the projected noise is one-dimensional Gaussian
$\langle N_{t},u_{i}\rangle\sim\mathcal{N}(0,v_{t,i}),\quad v_{t,i}=u_{i}^{\top}\Sigma_{t}u_{i}>0,$
and the two conditionals under $s_{i}$ and $s_{j}$ differ only by a mean shift along $u_{i}$. Define
$a(X,X^{\prime};R_{t})$ $=\langle\Delta_{t}(X,X^{\prime};R_{t}),u_{i}\rangle$
$=\langle f_{t}(X,y_{<t};R_{t})-f_{t}(X^{\prime},y_{<t};R_{t}),u_{i}\rangle.$
Then, conditioned on a realization $((X,X^{\prime}),R_{t})$, we can write
$\widetilde{Y}_{t}\mid(s_{i},\theta,X,X^{\prime},R_{t})$ $\sim\mathcal{N}(\mu+a,v_{t,i}),$
$\widetilde{Y}_{t}\mid(s_{j},\theta,X,X^{\prime},R_{t})$ $\sim\mathcal{N}(\mu,v_{t,i}),$
for some baseline mean $\mu$ that cancels in the divergence.
For two 1-D Gaussians with the same variance $v_{t,i}$ and mean difference $a$, the Rényi divergence of order $\alpha>1$ is
$\mathtt{D}_{\alpha}\bigl(\mathcal{N}(\mu+a,v_{t,i})\big\|\mathcal{N}(\mu,v_{t,i})\bigr)=\frac{\alpha a^{2}}{2v_{t,i}}.$ (39)
By the HUC bound (38), we have $|a(X,X^{\prime};R_{t})|\leq\sqrt{h_{t,i}}$ for $\gamma\times\mathbb{P}_{\eta,\rho}$-almost every $((X,X^{\prime}),R_{t})$, so
$\mathtt{D}_{\alpha}\bigl(\mathcal{N}(\mu+a,v_{t,i})\,\big\|\,\mathcal{N}(\mu,v_{t,i})\bigr)\leq\frac{\alpha}{2v_{t,i}}h_{t,i}.$ (40)
Step 2. Let $S:=(X,X^{\prime},R_{t})$ denote the joint randomness with probability measure $\nu:=\gamma\times\mathbb{P}_{\eta,\rho}$. For each realization $s\in\mathrm{supp}(\nu)$, define
$\mathrm{P}_{s}:=\mathcal{N}(\mu+a(s),v_{t,i}),\quad\mathrm{Q}_{s}:=\mathcal{N}(\mu,v_{t,i}).$
Then the 1-D marginals under secrets $s_{i}$ and $s_{j}$ can be written as mixtures with the same mixing measure $\nu$:
$\mathrm{P}:=\Pr\big(\langle Y_{t},u_{i}\rangle\mid s_{i},\theta\big)=\int\mathrm{P}_{s}\nu(ds),$
$\mathrm{Q}:=\Pr\big(\langle Y_{t},u_{i}\rangle\mid s_{j},\theta\big)=\int\mathrm{Q}_{s}\nu(ds).$
We now use the following standard inequality for Rényi divergence under mixtures with common weights: for $\alpha>1$ and any probability measure $\nu$,
$\mathtt{D}_{\alpha}\Bigl(\int\mathrm{P}_{s}\nu(ds)\,\Big\|\,\int\mathrm{Q}_{s}\nu(ds)\Bigr)\leq\operatorname*{ess\,sup}_{s}\mathtt{D}_{\alpha}(\mathrm{P}_{s}\|\mathrm{Q}_{s}).$ (41)
One way to see (41) is to consider the joint distributions on $(S,Z)$ under two secrets:
$R(ds,dz):=\nu(ds)\mathrm{P}_{s}(dz)\quad\textup{and}\quad\widetilde{R}(ds,dz):=\nu(ds)\mathrm{Q}_{s}(dz).$
By the data-processing inequality for Rényi divergence (applied to the channel $(S,Z)\mapsto Z$),
[…]
$\mathtt{D}_{\alpha}(R\|\widetilde{R})\leq\operatorname*{ess\,sup}_{s}\mathtt{D}_{\alpha}(\mathrm{P}_{s}\|\mathrm{Q}_{s}),$
which yields (41).
Applying (41) to $P_{s},Q_{s}$ and using the pointwise bound (40), we obtain
$\mathtt{D}_{\alpha}\left(\Pr\big(\langle Y_{t},u_{i}\rangle\mid s_{i},\theta\big)\,\Big\|\,\Pr\big(\langle Y_{t},u_{i}\rangle\mid s_{j},\theta\big)\right)$ $=\mathtt{D}_{\alpha}(\mathrm{P}\|\mathrm{Q})$
$\leq\operatorname*{ess\,sup}_{s}\mathtt{D}_{\alpha}(\mathrm{P}_{s}\|\mathrm{Q}_{s})$
$\leq\frac{\alpha}{2v_{t,i}}h_{t,i}.$
Since $\theta$, $(s_{i},s_{j})$, $y_{<t}$, and $\gamma$ were arbitrary, the bound
@@
$\mathtt{D}_{\alpha}\left(\Pr\big(\langle Y_{t},u_{i}\rangle\mid s_{i},\theta\big)\,\Big\|\,\Pr\big(\langle Y_{t},u_{i}\rangle\mid s_{j},\theta\big)\right)\leq\frac{\alpha}{2}\frac{h_{t,i}}{v_{t,i}}$
holds uniformly over all admissible choices. ∎
Appendix L Proof of Theorem 5.6
@@
Step 1. Fix $\alpha>1$, a prior $\theta\in\Theta$, and a secret pair $(s_{i},s_{j})\in\mathcal{Q}$ with positive prior mass. Fix an arbitrary history $y_{<t}$ in the support. For this fixed $y_{<t}$, the HUC condition (42) holds with the same vector $h_{t}$. Therefore, Lemma 5.1 applies to the conditional channel
[…]
with covariance $\Sigma_{t}=\sigma^{2}I_{d}$ and HUC vector $h_{t}$.
For each slice $u_{\ell}\in\mathcal{U}$, Lemma 5.1 together with $v_{t,\ell}:=u_{\ell}^{\top}\Sigma_{t}u_{\ell}=\sigma^{2}$ gives
$\mathtt{D}_{\alpha}\Bigl(\Pr\bigl(\langle Y_{t},u_{\ell}\rangle\mid s_{i},\theta,y_{<t}\bigr)\,\Big\|\,\Pr\bigl(\langle Y_{t},u_{\ell}\rangle\mid s_{j},\theta,y_{<t}\bigr)\Bigr)$ (43)
$\leq\frac{\alpha}{2}\frac{h_{t,\ell}}{v_{t,\ell}}=\frac{\alpha}{2\sigma^{2}}h_{t,\ell}.$
Define
$\varepsilon_{t,\ell}:=\frac{\alpha}{2\sigma^{2}}h_{t,\ell}.$
By (42), the bound (43) holds uniformly over all histories $y_{<t}$, all priors $\theta$, all $(s_{i},s_{j})\in\mathcal{Q}$, and all admissible couplings.
@@
$\mathtt{D}_{\alpha}\bigl(P^{(\ell)}_{t,y_{<t}}\,\big\|\,Q^{(\ell)}_{t,y_{<t}}\bigr)\leq\varepsilon_{t,\ell},\qquad\varepsilon_{t,\ell}:=\frac{\alpha}{2\sigma^{2}}\,h_{t,\ell}.$ (44)
@@
$\mathrm{P}^{(\ell)}:=\Pr\bigl(Z^{(\ell)}_{1:T}\mid s_{i},\theta\bigr),\qquad\mathrm{Q}^{(\ell)}:=\Pr\bigl(Z^{(\ell)}_{1:T}\mid s_{j},\theta\bigr)$
@@
The family $\{\mathrm{P}^{(\ell)}_{t,y_{<t}},\mathrm{Q}^{(\ell)}_{t,y_{<t}}\}_{t,y_{<t}}$ satisfies the assumptions of Lemma L.1 (stated later) with per–step bounds $\varepsilon_{t}=\varepsilon_{t,\ell}$. Applying Lemma L.1 yields
@@
$\mathtt{ASD}^{\omega}_{\alpha}\Bigl(\Pr(M(X)\mid s_{i},\theta)\,\Big\|\,\Pr(M(X)\mid s_{j},\theta)\Bigr)$
@@
$\mathtt{ASD}^{\omega}_{\alpha}\Bigl(\Pr(M(X)\mid s_{i},\theta)\,\Big\|\,\Pr(M(X)\mid s_{j},\theta)\Bigr)$ $\leq\sum_{\ell=1}^{m}\omega_{\ell}\frac{\alpha}{2\sigma^{2}}\sum_{t=1}^{T}h_{t,\ell}$
@@
then for every prior $\theta$ and every $(s_{i},s_{j})\in\mathcal{Q}$ with positive prior mass, the $\omega$–Ave-SRD between the corresponding output laws is at most $\epsilon$, i.e., Algorithm 2 is $(\alpha,\epsilon,\omega)$–Ave-SRPP.
Step 4. For Joint-SRPP, we use the joint sliced Rényi divergence (Definition 4.3), which for the two output laws can be written as
@@
where, for each $u_{\ell}\in\mathcal{U}$, $\mathrm{P}_{u_{\ell}}$ and $\mathrm{Q}_{u_{\ell}}$ denote the laws of $Z^{(\ell)}_{1:T}$ under $s_{i}$ and $s_{j}$, respectively, and $\mathrm{P}_{U},\mathrm{Q}_{U}$ are the corresponding random choices when $U\sim\omega$.
@@
then for every prior $\theta$ and every $(s_{i},s_{j})\in\mathcal{Q}$ the Joint-SRD between the corresponding output laws is at most $\epsilon$, i.e., Algorithm 2 is $(\alpha,\epsilon,\omega)$–Joint-SRPP. ∎
Lemma L.1 (Adaptive composition of Rényi divergence).
Let $\alpha>1$ and let $\mathrm{P}$ and $\mathrm{Q}$ be probability measures on the product space $\mathcal{Y}_{1}\times\cdots\times\mathcal{Y}_{T}$ with coordinate random variables $Y_{1},\dots,Y_{T}$. For each $t=1,\dots,T$, let $\mathrm{P}_{t\mid y_{<t}}$ and $\mathrm{Q}_{t\mid y_{<t}}$ denote regular conditional distributions of $Y_{t}$ given $Y_{<t}=y_{<t}$ under PP and $Q$, respectively, where $y_{<t}=(y_{1},\dots,y_{t-1})$. Suppose there exist constants $\varepsilon_{t}\geq 0$ such that
@@
$\mathtt{D}_{\alpha}(\mathrm{P}\,\|\,\mathrm{Q})=\frac{1}{\alpha-1}\log\mathbb{E}_{\mathrm{Q}}\bigl[L(Y_{1:T})^{\alpha}\bigr]=\frac{1}{\alpha-1}\log\mathbb{E}_{\mathrm{Q}}\Bigl[\prod_{t=1}^{T}L_{t}(Y_{\leq t})^{\alpha}\Bigr].$
@@
Appendix M Proof of Theorem 5.7
@@
Fix a slice profile $\{\mathcal{U},\omega\}$ with $\mathcal{U}=\{u_{\ell}\}_{\ell=1}^{m}\subset\mathbb{S}^{d-1}$ and $\omega\in\Delta(\mathcal{U})$. Let $\alpha>1$ and fix an arbitrary prior $\theta\in\Theta$ and secret pair $(s_{i},s_{j})\in\mathcal{Q}$.
[…]
For each iteration $t$ and slice $u_{\ell}\in\mathcal{U}$, define the (random) directional shift
$\widetilde{\Delta}_{t,\ell}(X,X^{\prime};\circledR):=\big\langle f_{t}(X,y_{<t};R_{t}(\circledR))-f_{t}(X^{\prime},y_{<t};R_{t}(\circledR)),u_{\ell}\big\rangle,$
[…]
For each realization $\circledR=\xi$, let
$h_{t,\ell}(\xi)=\operatorname*{ess\,sup}_{((X,X^{\prime}),R_{t})}\bigl|\Delta_{t,\ell}(X,X^{\prime};\xi)\bigr|^{2}.$
Then, for every fixed $\xi$, the vector $h_{t}(\xi)=(h_{t,1}(\xi),\dots,h_{t,m}(\xi))$ is a valid HUC vector at iteration $t$ in the sense of Definition 5.1: it uniformly bounds the squared directional shifts along each slice $u_{\ell}$ over all priors, secrets, couplings, and histories, with the subsampling pattern held fixed.
Step 1. Fix an arbitrary realization $\circledR=\xi$. By Lemma 5.1, with isotropic Gaussian noise $N_{t}\sim\mathcal{N}(0,\sigma^{2}I_{d})$, so that the projected variance is $v_{t,\ell}=u_{\ell}^{\top}\Sigma_{t}u_{\ell}=\sigma^{2}$ for all $t,\ell$, the per-iteration, per-slice Rényi cost satisfies
@@
$\mathtt{D}_{\alpha}\left(\Pr\big(\langle Y_{t},u_{\ell}\rangle\mid s_{i},\theta,\circledR=\xi\big)\;\Big\|\;\Pr\big(\langle Y_{t},u_{\ell}\rangle\mid s_{j},\theta,\circledR=\xi\big)\right)$
$\leq\frac{\alpha}{2}\frac{h_{t,\ell}(\xi)}{\sigma^{2}}.$
Applying the SRPP accountant of Theorem 5.6 to the conditional mechanism $\mathcal{M}^{\theta}_{\xi}$ (Algorithm 2 with subsampling fixed to $\xi$), we obtain
$\mathtt{AveSD}^{\omega}_{\alpha}\bigl(\mathcal{M}^{\theta}_{\xi}(s_{i})\,\|\,\mathcal{M}^{\theta}_{\xi}(s_{j})\bigr)\leq\frac{\alpha}{2\sigma^{2}}\sum_{t=1}^{T}\sum_{\ell=1}^{m}\omega_{\ell}h_{t,\ell}(\xi),$ (46)
and
$\mathtt{JSD}^{\omega}_{\alpha}\bigl(\mathcal{M}^{\theta}_{\xi}(s_{i})\,\|\,\mathcal{M}^{\theta}_{\xi}(s_{j})\bigr)\leq\frac{\alpha}{2\sigma^{2}}\sum_{t=1}^{T}\max_{\ell\in[m]}h_{t,\ell}(\xi).$ (47)
These inequalities hold for every fixed subsampling pattern $\xi$, every prior $\theta$, and every secret pair $(s_{i},s_{j})\in\mathcal{Q}$.

### v3

@@ Appendix H Minimality and Attainability of HUC
Fix $t$, a slice $u_{i}\in\mathbb{S}^{d-1}$, a prior $\theta\in\Theta$, and a secret pair $(s_{i},s_{j})\in\mathcal{Q}$ with $P^{S}_{\theta}(s_{i}),P^{S}_{\theta}(s_{j})>0$. Let $\gamma\in\Pi(\mu^{\theta}_{s_{i}},\mu^{\theta}_{s_{j}})$ and draw $((X,X^{\prime}),R_{t})\sim\gamma\times\mathbb{P}_{\eta,\rho}$. Let $N_{t}\sim\mathcal{N}(0,\Sigma_{t})$ be independent of $((X,X^{\prime}),R_{t})$ and set
$\widetilde{Y}_{t}=\langle f_{t}(X,y_{<t};R_{t}),u_{i}\rangle+\langle N_{t},u_{i}\rangle,$
$\widetilde{Y}_{t}^{\prime}=\langle f_{t}(X^{\prime},y_{<t};R_{t}),u_{i}\rangle+\langle N_{t},u_{i}\rangle.$
Then $\langle N_{t},u_{i}\rangle\sim\mathcal{N}(0,v_{t,i})$ with $v_{t,i}=u_{i}^{\top}\Sigma_{t}u_{i}>0$.
Step 1. Let $S:=(X,X^{\prime},R_{t})$ and define the (random) mean shift
$a(S):=\langle f_{t}(X,y_{<t};R_{t})-f_{t}(X^{\prime},y_{<t};R_{t}),u_{i}\rangle.$
Conditioned on $S$, we have
$\widetilde{Y}_{t}\mid S\sim\mathcal{N}(m(S)+a(S),\,v_{t,i}),\quad\widetilde{Y}_{t}^{\prime}\mid S\sim\mathcal{N}(m(S),\,v_{t,i}),$
for some $m(S)\in\mathbb{R}$ (which cancels in the divergence). Hence for every realization $S$,
$\mathtt{D}_{\alpha}\!\left(\widetilde{Y}_{t}\mid S\ \big\|\ \widetilde{Y}_{t}^{\prime}\mid S\right)$ $=\mathtt{D}_{\alpha}\!\left(\mathcal{N}(m+a,v_{t,i})\,\big\|\,\mathcal{N}(m,v_{t,i})\right)$
$=\frac{\alpha\,a(S)^{2}}{2v_{t,i}}.$
By the HUC property (Definition 5.1), we have $a(S)^{2}\leq h_{t,i}$ $\gamma\times\mathbb{P}_{\eta,\rho}$-a.s., therefore
$\mathtt{D}_{\alpha}\!\left(\widetilde{Y}_{t}\mid S\ \big\|\ \widetilde{Y}_{t}^{\prime}\mid S\right)\leq\frac{\alpha}{2v_{t,i}}\,h_{t,i}\quad\gamma\times\mathbb{P}_{\eta,\rho}\text{-a.s.}$
Step 2. Let $\nu$ be the law of $S$ under $\gamma\times\mathbb{P}_{\eta,\rho}$. For each $s\in\mathrm{supp}(\nu)$, write $P_{s}$ as the probability distribution of $\widetilde{Y}_{t}$ given $s$ and $Q_{s}$ as the probability distribution of $\widetilde{Y}^{\prime}_{t}$ given $s$. Then the unconditional one-dimensional output laws can be written as mixtures with the same weights:
$P=\Pr(\langle Y_{t},u_{i}\rangle\mid s_{i},\theta)=\int P_{s}\,\nu(ds),$
$Q=\Pr(\langle Y_{t},u_{i}\rangle\mid s_{j},\theta)=\int Q_{s}\,\nu(ds).$
By the standard Rényi-divergence bound for mixtures with common weights (equivalently, by applying data processing to the joint laws $\nu(ds)P_{s}(dy)$ and $\nu(ds)Q_{s}(dy)$ and then marginalizing out $S$), for any $\alpha>1$,
$\mathtt{D}_{\alpha}(P\|Q)\leq\operatorname*{ess\,sup}_{s}\mathtt{D}_{\alpha}(P_{s}\|Q_{s}).$
Combining with the bound from Step 1 yields
@@
$\mathtt{D}_{\alpha}\!\left(\Pr(\langle Y_{t},u_{i}\rangle\mid s_{i},\theta)\ \big\|\ \Pr(\langle Y_{t},u_{i}\rangle\mid s_{j},\theta)\right)\leq\frac{\alpha}{2}\frac{h_{t,i}}{v_{t,i}}.$
Since $\theta$ and $(s_{i},s_{j})$ were arbitrary, the claim follows. ∎
Appendix T Proof of Theorem 5.6
@@
Step 1. Fix $\alpha>1$, $\theta\in\Theta$, and $(s_{i},s_{j})\in\mathcal{Q}$ with positive prior mass. Fix an arbitrary history $y_{<t}$ in the support. Fix an arbitrary coupling $\gamma\in\Pi(\mu^{\theta}_{s_{i}},\mu^{\theta}_{s_{j}})$.
Since $h_{t}$ is a valid HUC vector, for $\gamma\times\mathbb{P}_{\eta,\rho}$-a.s. in $((X,X^{\prime}),R_{t})$,
$\bigl|\langle f_{t}(X,y_{<t};R_{t})-f_{t}(X^{\prime},y_{<t};R_{t}),u_{\ell}\rangle\bigr|^{2}\leq h_{t,\ell},\quad\forall\ell\in[m].$
By Fubini/Tonelli, there exists a measurable set $\mathcal{R}_{t}$ with $\Pr(R_{t}\in\mathcal{R}_{t})=1$ such that for every $r\in\mathcal{R}_{t}$ the same bound holds $\gamma$-a.s. in $(X,X^{\prime})$:
$\bigl|\langle f_{t}(X,y_{<t};r)-f_{t}(X^{\prime},y_{<t};r),u_{\ell}\rangle\bigr|^{2}\leq h_{t,\ell}.$
Condition on $R_{t}=r\in\mathcal{R}_{t}$. Then $\langle Y_{t},u_{\ell}\rangle$ is Gaussian with variance $v_{t,\ell}=u_{\ell}^{\top}(\sigma^{2}I_{d})u_{\ell}=\sigma^{2}$ under both secrets, and the mean shift is $\langle f_{t}(X,y_{<t};r)-f_{t}(X^{\prime},y_{<t};r),u_{\ell}\rangle$. Hence Lemma 5.1 yields, for all $r\in\mathcal{R}_{t}$,
$\mathtt{D}_{\alpha}\!\Big($ $\Pr(\langle Y_{t},u_{\ell}\rangle\mid s_{i},\theta,y_{<t},R_{t}=r)$ (68)
$\ \Big\|\ \Pr(\langle Y_{t},u_{\ell}\rangle\mid s_{j},\theta,y_{<t},R_{t}=r)\Big)\leq\frac{\alpha}{2\sigma^{2}}\,h_{t,\ell}.$
Now view $(R_{t},\langle Y_{t},u_{\ell}\rangle)$ as the released pair at round $t$. Because $R_{t}$ is independent of the secret (and of $y_{<t}$), it has the same conditional law under $s_{i}$ and $s_{j}$. Using (68) (which holds for $R_{t}$-a.e. $r$) and the definition of Rényi divergence, we obtain
$\mathtt{D}_{\alpha}\!\Big(\Pr(R_{t},\langle Y_{t},u_{\ell}\rangle\mid s_{i},\theta,y_{<t})\ \Big\|\ \Pr(R_{t},\langle Y_{t},u_{\ell}\rangle\mid s_{j},\theta,y_{<t})\Big)$
$\leq\frac{\alpha}{2\sigma^{2}}\,h_{t,\ell}.$
Finally, marginalizing out $R_{t}$ is post-processing, so by data processing,
$\mathtt{D}_{\alpha}\!\Big(\Pr(\langle Y_{t},u_{\ell}\rangle\mid s_{i},\theta,y_{<t})\ \Big\|\ \Pr(\langle Y_{t},u_{\ell}\rangle\mid s_{j},\theta,y_{<t})\Big)\leq\frac{\alpha}{2\sigma^{2}}\,h_{t,\ell}.$ (69)
This bound holds uniformly over all histories $y_{<t}$ in the support.
@@
$\mathtt{D}_{\alpha}\bigl(P^{(\ell)}_{t,y_{<t}}\,\big\|\,Q^{(\ell)}_{t,y_{<t}}\bigr)\leq\varepsilon_{t,\ell},\quad\varepsilon_{t,\ell}:=\frac{\alpha}{2\sigma^{2}}\,h_{t,\ell}.$ (70)
@@
$\mathrm{P}^{(\ell)}:=\Pr\bigl(Z^{(\ell)}_{1:T}\mid s_{i},\theta\bigr),\quad\mathrm{Q}^{(\ell)}:=\Pr\bigl(Z^{(\ell)}_{1:T}\mid s_{j},\theta\bigr)$
@@
The family $\{\mathrm{P}^{(\ell)}_{t,y_{<t}},\mathrm{Q}^{(\ell)}_{t,y_{<t}}\}_{t,y_{<t}}$ satisfies the assumptions of Lemma T.1 (stated later) with per–step bounds $\varepsilon_{t}=\varepsilon_{t,\ell}$. Applying Lemma T.1 yields
@@
$\mathtt{AveSD}^{\omega}_{\alpha}\Bigl(\Pr(M(X)\mid s_{i},\theta)\,\Big\|\,\Pr(M(X)\mid s_{j},\theta)\Bigr)$
@@
$\mathtt{AveSD}^{\omega}_{\alpha}\Bigl(\Pr(M(X)\mid s_{i},\theta)\,\Big\|\,\Pr(M(X)\mid s_{j},\theta)\Bigr)$
$\leq\sum_{\ell=1}^{m}\omega_{\ell}\frac{\alpha}{2\sigma^{2}}\sum_{t=1}^{T}h_{t,\ell}$
@@
then for every prior $\theta$ and every $(s_{i},s_{j})\in\mathcal{Q}$ with positive prior mass, the $\omega$–Ave-SRD between the corresponding output distributions is at most $\epsilon$, i.e., Algorithm 1 is $(\alpha,\epsilon,\omega)$–Ave-SRPP.
Step 4. For Joint-SRPP, we use the joint sliced Rényi divergence (Definition 4.3), which for the two output distributions can be written as
@@
where, for each $u_{\ell}\in\mathcal{U}$, $\mathrm{P}_{u_{\ell}}$ and $\mathrm{Q}_{u_{\ell}}$ denote the distributions of $Z^{(\ell)}_{1:T}$ under $s_{i}$ and $s_{j}$, respectively, and $\mathrm{P}_{U},\mathrm{Q}_{U}$ are the corresponding random choices when $U\sim\omega$.
@@
then for every prior $\theta$ and every $(s_{i},s_{j})\in\mathcal{Q}$ the Joint-SRD between the corresponding output distributions is at most $\epsilon$, i.e., Algorithm 1 is $(\alpha,\epsilon,\omega)$–Joint-SRPP. ∎
Lemma T.1 (Adaptive composition of Rényi divergence).
Let $\alpha>1$ and let $\mathrm{P}$ and $\mathrm{Q}$ be probability measures on the product space $\mathcal{Y}_{1}\times\cdots\times\mathcal{Y}_{T}$ with coordinate random variables $Y_{1},\dots,Y_{T}$. For each $t=1,\dots,T$, let $\mathrm{P}_{t\mid y_{<t}}$ and $\mathrm{Q}_{t\mid y_{<t}}$ denote regular conditional distributions of $Y_{t}$ given $Y_{<t}=y_{<t}$ under $P$ and $Q$, respectively, where $y_{<t}=(y_{1},\dots,y_{t-1})$. Suppose there exist constants $\varepsilon_{t}\geq 0$ such that
@@
$\mathtt{D}_{\alpha}(\mathrm{P}\,\|\,\mathrm{Q})=\frac{1}{\alpha-1}\log\mathbb{E}_{\mathrm{Q}}\bigl[L(Y_{1:T})^{\alpha}\bigr]$
$=\frac{1}{\alpha-1}\log\mathbb{E}_{\mathrm{Q}}\Bigl[\prod_{t=1}^{T}L_{t}(Y_{\leq t})^{\alpha}\Bigr].$
@@
Appendix U Proof of Theorem 5.7
@@
Fix a slice profile $(\mathcal{U},\omega)$ with $\mathcal{U}=\{u_{\ell}\}_{\ell=1}^{m}$ and $\omega\in\Delta(\mathcal{U})$. Fix $\alpha>1$. Fix an arbitrary prior $\theta\in\Theta$ and an arbitrary secret pair $(s_{i},s_{j})\in\mathcal{Q}$.
Let $\eta=(\eta_{1},\dots,\eta_{T})$ denote the (possibly time-varying) subsampling parameters used by Algorithm 1 (e.g., sampling rates). Let $R_{t}$ denote the realized subsampling outcome at round $t$ (e.g., the sampled index multiset), and let $\mathbf{R}:=(R_{1},\dots,R_{T})$. For each realization $\mathbf{r}$ of $\mathbf{R}$, let $\mathcal{M}^{\theta,\mathbf{r}}_{s}$ denote the conditional law of the final output of Algorithm 1 under prior $\theta$ and secret $s$, conditional on $\mathbf{R}=\mathbf{r}$.
Throughout, the Gaussian noises are i.i.d. $N_{t}\sim\mathcal{N}(0,\sigma^{2}I_{d})$, independent of all data and of $\mathbf{R}$. Hence for every slice $u_{\ell}$,
$v_{t,\ell}:=u_{\ell}^{\top}(\sigma^{2}I_{d})u_{\ell}=\sigma^{2}.$
Lemma U.1 (Expected Gaussian sliced Rényi cost).
Fix iteration $t$ and direction $u\in\mathbb{S}^{d-1}$. Let $N_{t}\sim\mathcal{N}(0,\sigma^{2}I_{d})$ be independent of all data and of $R_{t}$. Fix $\theta\in\Theta$, $(s_{i},s_{j})\in\mathcal{Q}$, a history $y_{<t}$ in the support, and a coupling $\gamma\in\Pi(\mu_{s_{i}}^{\theta},\mu_{s_{j}}^{\theta})$. For $\gamma$-a.e. $(X,X^{\prime})$, define the random directional shift
$\Delta_{t}^{u}(X,X^{\prime};R_{t}):=\big\langle f_{t}(X,y_{<t};R_{t})-f_{t}(X^{\prime},y_{<t};R_{t}),u\big\rangle.$
Then, for any $\alpha>1$,
@@
$\mathbb{E}_{R_{t}}\!\left[\mathtt{D}_{\alpha}\!\left(\Pr(\langle Y_{t},u\rangle\mid s_{i},\theta,R_{t})\ \Big\|\ \Pr(\langle Y_{t},u\rangle\mid s_{j},\theta,R_{t})\right)\right]$
$=\frac{\alpha}{2\sigma^{2}}\,\mathbb{E}_{R_{t}}\!\left[\big|\Delta_{t}^{u}(X,X^{\prime};R_{t})\big|^{2}\right],$
where $Y_{t}=f_{t}(X,y_{<t};R_{t})+N_{t}$.
Proof.
Fix $t,u,\theta,(s_{i},s_{j}),y_{<t},\gamma$ as stated and take $\gamma$-a.e. $(X,X^{\prime})$. Condition on $R_{t}=r$. Then $\langle Y_{t},u\rangle$ is Gaussian with variance $\sigma^{2}$ and mean $\langle f_{t}(X,y_{<t};r),u\rangle$, while the corresponding mean under secret $s_{j}$ is $\langle f_{t}(X^{\prime},y_{<t};r),u\rangle$. Hence the conditional mean shift equals $\Delta_{t}^{u}(X,X^{\prime};r)$ and the standard Gaussian identity gives
$\mathtt{D}_{\alpha}\!\left(\Pr(\langle Y_{t},u\rangle\mid s_{i},\theta,R_{t}=r)\ \Big\|\ \Pr(\langle Y_{t},u\rangle\mid s_{j},\theta,R_{t}=r)\right)$
$=\frac{\alpha}{2\sigma^{2}}\,\big|\Delta_{t}^{u}(X,X^{\prime};r)\big|^{2}.$
Taking expectation over $R_{t}$ yields the claim. ∎